OpenStudy (anonymous):

Given that log 9 y=x, show that log y 3=1/2x.

OpenStudy (cwrw238):

does 1/2x mean 1 divided by 2x, or half x?

OpenStudy (anonymous):

half, like 1 over 2x

jhonyy9 (jhonyy9):

- so do you know how you can changing the base of one logarithm on one different what is added ? like for example : loga(b) = ? in base of b ,like logb(a)

jhonyy9 (jhonyy9):

- because log9(y)=x - and you need to show that logy(3) = 1/(2x)

jhonyy9 (jhonyy9):

- right ?

OpenStudy (anonymous):

- i know how to change the base - and right.

jhonyy9 (jhonyy9):

yes and after that you need to write log9(y) in the form of log y (9) - and from this you will ned to see that log y (3) = 1/2x

OpenStudy (anonymous):

wait, i'm a bit slow.

jhonyy9 (jhonyy9):

ok

OpenStudy (anonymous):

how about you show the steps?

jhonyy9 (jhonyy9):

how you write log9(y) in the form of log y(3) ?

jhonyy9 (jhonyy9):

- and can you tel me please that there is logy(3) = (1/2) x or 1/(2x) ?

OpenStudy (anonymous):

Argh, now i don't know anything anymore. The question stated logy 3= 1/(2x)

jhonyy9 (jhonyy9):

- so one second possibile step if you begin it from logy (3) =1/(2x) and here change x by log9 (y) so than will get logy(3)= 1/(2log9 (y) - but now to continue you need changing the base of 2log9 (y) on y

OpenStudy (anonymous):

- okay, i got the first one. - to change the base of 2log9 (y) on y....how?

OpenStudy (anonymous):

\[\log_{9}y=x \]\[\log_{y}3=1/(2x) \]Change the base of \[\log_{y}3=1/(2x) \]\[\log_{9}3/\log_{9}y \]Substitute \[\log_{9}y=x \]\[\log_{9}3/x=1/(2x) \]\[\log_{9}3=x/(2x) \]\[\log_{9}3=1/2 \]\[9^{(1/2)}=3\]